OpenStudy (anonymous):

Qudratic equations..x^2 - 4x = 12

OpenStudy (anonymous):

imaginary roots.no real roots

OpenStudy (anonymous):

\[x^2-4x-12=0\] \[(x-6)(x+2)=0\] so \[x=6\ or \ x=-2\]

OpenStudy (anonymous):

Ok that what I got but the answer sheet is telling me 2,6,12,16

OpenStudy (anonymous):

solve it using completing the square method add square of half of the coefficient of x i.e. 4/2 = 2 to both sides x² - 4x + 2² = 12 + 2² LHS takes the form a² - 2ab + b² = (a - b)² (x - 2)² = 16 taking root of both sides x-2 = ± root(16) x = 2 ± 4 x = 2+4 =6 and x = 2-4 = -2

OpenStudy (anonymous):

x = 6, x = -2

OpenStudy (anonymous):

The answer sheet is wrong.

OpenStudy (anonymous):

6

OpenStudy (anonymous):

or -2

OpenStudy (anonymous):

u can substitute the values into the equation and calculate... you will that -2 and 6 r the right answers..